OpenStudy (anonymous):

How do I polar plot the function r=2sin(30)sin(theta) 0<(theta)<2pi?

OpenStudy (anonymous):

r=sin(30)sin(theta) ?

OpenStudy (anonymous):

yes that is correct

jhonyy9 (jhonyy9):

but than you know that sin 30 = 1/2 so than will result r=sin(theta) yes ?

OpenStudy (gorv):

\[r=2*\frac{ 1 }{ 2 }*\sin \theta\]

OpenStudy (gorv):

r=sin (theta )

OpenStudy (gorv):

it is simply the sin